OpenStudy (anonymous):

what is the quotient?? (x^3 +7x^2 +5x-20) / (x+3) ... I know you would use synthetic division but i'm confused

OpenStudy (anonymous):

you can use either synthethic or long division

OpenStudy (anonymous):

1 7 5 -20 -3 _____________________________ 1

OpenStudy (anonymous):

-3 * 1 = -3 1 7 5 -20 -3 -3 _____________________________ 1

OpenStudy (anonymous):

7-3=4 1 7 5 -20 -3 -3 _____________________________ 1 4

OpenStudy (anonymous):

-3*4=-12 1 7 5 -20 -3 -3 -12 _____________________________ 1 4

OpenStudy (anonymous):

5-12=-7 1 7 5 -20 -3 -3 -12 _____________________________ 1 4 -7

OpenStudy (anonymous):

-3*-7=21 1 7 5 -20 -3 -3 -12 21 _____________________________ 1 4 -7

OpenStudy (anonymous):

21-20=1 1 7 5 -20 -3 -3 -12 21 _____________________________ 1 4 -7 1

OpenStudy (anonymous):

divide a polynomial of degree 3 by a poly of degree 1 get a poly of degree 2. your answer is \[x^2+4x-7+\frac{1}{x+3}\]

OpenStudy (anonymous):

thank you !!!!!!!!!!!!1